myabell:

I need help I solve it and got 4/3 but I'm not sure work is below

myabell:

Extrinix:

So the equation of a prism is \(V=Bh\). Commonly known as \(V=lwh\).

Extrinix:

Now, looking at this you can see that there are 3 cubes by 4 cubes, all measuring \(\dfrac{1}{3}\), which would give us, \(V=\dfrac{3}{3}(\dfrac{4}{3})(\dfrac{3}{3})\) So we would be left with, \(V=1\dfrac{1}{3}\) Which means that \(\dfrac{4}{3}\) would be correct, good job.

myabell:

Can someone help me with this...

myabell:

Can someone help^^

Extrinix:

Ok so, the surface area of a pyramid equation is: \(SA=B+\dfrac{1}{2}Pl\)

Extrinix:

Given this, what is the equation when you add the inputs?